A 0.10 M solution of formic acid (HCOOH) is prepared in various solvents:

a. pure water

b. 0.10 M NaHCOO solution,

c. 0.10 M Ca(HCOO)2 solution, and

d. 0.10 M NaCl.

Compute the pH of each solution. The Ka of formic acid is 1.78 x 10



1
Expert's answer
2021-03-15T09:09:47-0400

"a)pH=1\/2(pK_a-log(c_0))=2.38pH"


"b)pH=pK_a+log\left(\frac{c(NaHCOO)}{c(HCOOH)}\right)=pK_a=3.75pH"


"c) pH=pK_a+log\left(\frac{2c(Ca(HCOO)_2)}{c(HCOOH)}\right)=pK_a+log(2)=4.05pH"


"d) pH =2.38"